On January 1, 2021, Allied Industries leased a high-performance conveyer to Karrier Company for a four-year period ending DecNeed help on requirement 2, journal entries 2 and 3.
Answer #1

Journal entries:

No Date General Journal Debit Credit
2 Amortization expenses A/c $ 185,434
Right-of-Use asset A/c $ 185,434
3 Cash A/c $ 204,000
Interest Revenue A/c $    58,952
Lease receivable A/c $ 145,048

Working:

2.

a Periodic Lease Amount $        204,000
b PVIFA 6%, 4 years 3.46511
c=a*b PV of periodic lease amount $        706,882
d Excess lease (348,000-304000) $          44,000
e PVF 6%, 4 Years 0.79209
f=d*e PV of excess lease $          34,852
g=c+f PV of lease payment $        741,734
h=g/4 Amortization expenses $        185,434 (2)
i Guaranteed residual value $        348,000
j=i*e PV of guaranteed residual value $        275,647
k=c+j PV of lessor payment $        982,530
l=k*6% Interest revenue $          58,952 [3]
m Cash $        204,000 [3]
n=m-l Lease receivable $        145,048 [3]